OpenStudy (wonderwomen):

what is the missing fraction 1/5-?=3/40

OpenStudy (anonymous):

the two fraction together would result in a denominator 40. Normally if you add or detract two fractions with different denominators you multiply the denominators to create a new one. So if 5.? = 40 then ?=40/5. What is ?

OpenStudy (wonderwomen):

2/5?

OpenStudy (anonymous):

If you divide 40 by 5, what do you get?

OpenStudy (wonderwomen):

9?

OpenStudy (anonymous):

No, 8

OpenStudy (anonymous):

45/5=9

OpenStudy (wonderwomen):

so it will be 1/8?

OpenStudy (anonymous):

Ok, so now you know that \[\frac{ 1 }{ 5 }+\frac{ ? }{ 8 }=\frac{ 3 }{ 40 }\] Now let's put both of those fractions to 40 \[\frac{ 1.8 }{ 5.8 }-\frac{ ? }{ 5.8 }=\frac{ 3 }{ 40 }\] \[\frac{ 8 }{ 40 }-\frac{ ? }{ 40 }=\frac{ 3 }{ 40 }\] That would mean the missing fraction is \[\frac{ 5 }{ 40 }\] Now let's simplify that back to /8

OpenStudy (anonymous):

And yes 5/40=1/8

OpenStudy (wonderwomen):

thank you

OpenStudy (anonymous):

Gladly done :-)